OpenStudy (anonymous):

Which pairs are solutions to the equation? 5xy + 9 = 44 (1, 7) and (4, 11) (1, 7) and (7, 1) (7, 1) and (2, 3) (4, 11) and (2, 3)

pooja195 (pooja195):

Plug in each of your answer choices in the equation see what coordinates make the equation true :)

pooja195 (pooja195):

A. 5(1)(7) + 9 = 44 5(4)(11) + 9 = 44 etc

OpenStudy (anonymous):

5x7+9=44 5xy=44-9 xy=35/5 xy=7

OpenStudy (anonymous):

x*y=7 1*7 or 7*1

OpenStudy (anonymous):

so the answer is B

pooja195 (pooja195):

That makes little to no sense. Its better to let the asker do the work